Harris Company manufactures and sells a single product.
1.

A partially completed schedule of the company’s total and per unit costs over the relevant range of 53,000 to 93,000 units produced and sold annually is given below: Complete the schedule of the company’s total and unit costs. (Round the variable cost and fixed cost to 2 decimal places.)

Units Produced & Sold
53000 73000 93000
Total Costs
Variable Costs $180,200 ? ?
Fixed Costs 390,000 ? ?
Total Costs $570,200 ? ?
Cost per Unit
Variable Cost ? ? ?
Fixed Cost ? ? ?
Total Unit per Cost ? ? ?
2.

Assume that the company produces and sells 83,000 units during the year at a selling price of $8.97 per unit. Prepare a contribution format income statement for the year.

Harris Company
Contribution Format Income Statement
Sales ?
Variable Expenses ?
Contribution Margin ?
Fixed Expense ?
Net Operating Income ?
Answer #1

Schedule of the company’s total and unit costs.

Units Produced & Sold
53000 73000 93000
Total Costs
Variable Costs $180,200 $248,200 $316,200
Fixed Costs 390,000 390,000 390,000
Total Costs $570,200 $638,200 $706,200
Cost per Unit
Variable Cost 3.4 3.4 3.4
Fixed Cost 7.36 5.34 4.19
Total Unit per Cost 10.76 8.74 7.59

Variable cost per unit = Total variable cost/Number of units produced

= 180,200/53,000

= $3.4

Total variable cost at 73,000 units = 73,000 x 3.4

= $248,200

Total variable cost at 93,000 units = 93,000 x 3.4

= $316,200

Fixed cost per unit = Total fixed cost/Number of units produced

Fixed cost per unit at 53,000 units = 390,000/53,000

= $7.36

Fixed cost per unit at 73,000 units = 390,000/73,000

= $5.34

Fixed cost per unit at 93,000 units = 390,000/93,000

= $4.19

2.

contribution format income statement

Harris Company
Contribution Format Income Statement
Sales 744,510
Variable Expenses (282,200)
Contribution Margin 462,310
Fixed Expense (390,000)
Net Operating Income 72,310